OpenStudy (anonymous):

Find the altitude of an equilateral triangle if one of its sides measures 12 cm.

OpenStudy (anonymous):

Height: \(\frac{\sqrt{3}}2 \times \)side

OpenStudy (anonymous):

\[6 \sqrt{3} \]

jhonyy9 (jhonyy9):

144=36 +h^2 h=sqrt(144-36) h=sqrt108 = or sin60=h/12 sqrt3 h ----- = ----- ---- h= 6sqrt3 2 12

jhonyy9 (jhonyy9):

my pleasure yw good luck bye

OpenStudy (callisto):

 Draw this diagram may help you find out the answer! Good luck :) PS: you can use sine ratio since it is a right angle triangle!~